St corporation lost its inventory in a fire and they have asked you to calculate the value of Ending Inventory based on the following info: Beginning inventory $250,000 , Sales $980,000 , Purchases $590.000, Sales returns $55,000 , Purchase returns $30,000 . Average Gross Profit margin 55% . Show All work!
Answer #1

Net sales = sales - sales returns

=980000-55000

= $ 925,000

Cost of goods sold =net sales - Gross profit margin

= 925,000 - 55%

=$ 416,250

Net purchases = purchases - purchase returns

= 590000-30000

= $ 560,000

Closing stock = (Opening stock + Net purchases ) - COGS

= (250,000 + 560,000 ) - 416250

= $ 393,750

If you are benefited from the solution then please LIKE ,

if disliked then please specify the reason .